### v1

# Fork, Explore, Commit: OS Primitives for Agentic Exploration

Abstract.
[…]
Agentic exploration, AI agents, Filesystem branching, Copy-on-write, FUSE, Operating system primitives
copyright: noneconference: Agentic OS Workshop, ASPLOS 2026; March 23, 2026; Pittsburgh, USA
1. Introduction
Modern AI agents autonomously execute multi-step tasks across domains such as software engineering (Yang et al., 2024a; Wang et al., 2025; Anthropic, 2025), computer use (Wang et al., 2024), and OS tuning (Zheng et al., 2024) through a reason-then-act loop (Yao et al., 2023b), where each action (shell commands, file edits, package installations) may produce irreversible side effects. Increasingly, agent systems explore multiple solution paths in parallel (Yao et al., 2023a; Shinn et al., 2023; Yang et al., 2024a): for example, a software engineering agent may try several candidate fixes simultaneously, committing only the one that passes tests. We term this pattern agentic exploration: the autonomous, parallel pursuit of multiple solution paths with selective commitment of results. It requires isolating two types of state across parallel branches: filesystem state (workspace modifications) and process state (execution context), with the ability to commit successful branches and discard failed ones.
Existing OS mechanisms do not meet these requirements individually, and composing them is fragile: filesystem branching primitives lack nested branching, commit semantics, and unprivileged operation; process management mechanisms lack reliable branch-aware termination and sibling isolation; and stitching these together in userspace introduces race windows between steps, error-prone cleanup on partial failure, and ordering dependencies that defeat correctness (Section 2). In practice, current agent frameworks resort to ad hoc solutions such as git stashing, temporary directories, or container clones (Chase, 2022; Richards, 2023), which incur significant overhead and cannot capture all filesystem modifications.
We introduce the branch context, a new OS abstraction that captures these requirements (Section 3). A branch context is an isolated execution environment, with a copy-on-write filesystem view and a process group, that follows a fork/explore/commit lifecycle with first-commit-wins resolution and support for nesting. We realize branch contexts in Linux through two complementary components:
  1. (1)
    BranchFS: A FUSE-based (Szeredi and others, 2026) filesystem that gives each branch context an isolated copy-on-write workspace, with O(1) creation, atomic commit to the parent, and automatic sibling invalidation, all without root privileges.
  2. (2)
    branch(): A proposed Linux syscall that spawns processes into branch contexts with reliable termination, kernel-enforced sibling isolation, and first-commit-wins coordination.
[…]
We make the following contributions:
  • The branch context, a new OS abstraction for agentic exploration with copy-on-write isolation, fork/explore/commit lifecycle, first-commit-wins resolution, and nesting (Section 3).
  • Two complementary Linux implementations: BranchFS, a FUSE-based branching filesystem (Section 4); and branch(), a proposed syscall for process coordination (Section 5).
  • A preliminary evaluation of BranchFS (Section 6).
2. Motivation
2.1. AI Agent Execution Patterns
Several agent patterns amplify the side-effect challenge introduced above by exploring multiple execution paths: Tree/Graph-of-Thoughts reasoning (Yao et al., 2023a; Besta et al., 2024) searches over candidate solutions in parallel, Reflexion (Shinn et al., 2023) retries tasks after self-critique requiring rollback of prior side effects, and software engineering agents such as SWE-agent (Yang et al., 2024a) and OpenHands (Wang et al., 2025) attempt multiple fixes each producing substantial filesystem changes. These patterns share a common need for lightweight, isolated workspaces. Current frameworks use ad hoc solutions: LangChain (Chase, 2022) and AutoGPT (Richards, 2023) rely on git stashing, temporary directories, or container clones, while Claude Code (Anthropic, 2025) provides per-file snapshots that cannot capture changes from shell commands (e.g., npm install) and do not support parallel branching. None adequately meets these needs.
2.2. Requirements of Agentic Exploration
The agent patterns above impose six requirements on the execution environment:
R1: Isolated parallel execution.
Because multiple exploration paths run simultaneously and may modify the same files (e.g., editing main.py or running npm install), each path needs its own isolated view of filesystem and process state. Without isolation, concurrent modifications corrupt shared state.
R2: Atomic commit with single-winner resolution.
When one path succeeds (e.g., tests pass), its changes must be applied atomically to the shared workspace. All sibling paths must be automatically invalidated, since they now operate on stale state and could silently corrupt the committed result. Failed paths must be discardable with no side effects.
R3: Hierarchical nesting.
Tree-of-Thoughts and similar patterns explore hierarchically: an agent tries three strategies, each of which may try sub-variants. The execution environment must support nested branches, where each level commits to its immediate parent.
R4: Complete filesystem coverage.
Agent actions go beyond editing source files: they run shell commands that create build artifacts, install packages, and modify dotfiles. The isolation mechanism must capture all filesystem modifications, not just tracked files.
R5: Lightweight, unprivileged, and portable.
Agents create and discard branches frequently (e.g., per LLM reasoning step). Branch creation must be sub-millisecond, require no root privileges, and work portably across underlying filesystems (ext4, XFS, NFS, etc.) since agents run in diverse environments. This rules out heavyweight mechanisms such as VM snapshots, privileged container runtimes, or filesystem-specific solutions.
R6: Process coordination.
Agents spawn processes (compilers, test runners, package managers) within each path. On commit or abort, all processes in a branch must be reliably terminated; sibling branches must be isolated from each other to prevent cross-path interference (e.g., one branch’s test runner killing another’s).
2.3. Limitations of Existing Mechanisms
No existing OS mechanism satisfies these requirements (Tables 1 and 2).
Filesystem branching.
Union filesystems such as UnionFS (Wright et al., 2006) and its successor OverlayFS (Brown, 2016) provide a single writable upper directory with no native commit-to-parent or sibling invalidation, and mounting requires root privileges. Btrfs (Btrfs Project, 2023) and ZFS (Bonwick et al., 2003) support nested subvolumes and clones but are each tied to a single filesystem type and lack native commit-to-parent. Device-mapper snapshots (Linux Kernel Documentation, 2023) require raw block devices with O(depth) read latency. DAXFS (Cong Wang, 2026) provides branching with commit and sibling invalidation but targets memory-backed storage and is not portable.
At coarser granularity, VM snapshot systems (QEMU Project, 2024; VMware, 2021; Cully et al., 2008; Lagar-Cavilla et al., 2009) have whole-VM overhead; recent serverless snapshot work (Agache et al., 2020; Ustiugov et al., 2021; Du et al., 2020) reduces startup latency but does not address branching semantics; and CRIU (CRIU Project, 2024) requires full state serialization, all violating R5.
Process management.
Process groups and sessions (Stevens and Rago, 2013) allow group termination, but processes can escape via setsid() or setpgid(); manual PID tracking via /proc suffers from race conditions. Cgroups (Heo, 2015) provide reliable termination but require setup and typically need root privileges (cgroup v2 delegation via systemd can avoid root but requires prior configuration), and do not prevent signaling between siblings. PID namespaces provide complete isolation but impose PID 1 init overhead. The clone() syscall (Kerrisk, 2024) enables fine-grained resource sharing but requires combining multiple mechanisms. None satisfies R6 in full.
Table 1. Feature comparison of filesystem branching mechanisms.
Feature OverlayFS/UnionFS Btrfs/ZFS DM-Snap DAXFS BranchFS
Portable across FSes
Nested branches
Commit-to-parent
Sibling invalidation
No root privileges
Table 2. Feature comparison of process management mechanisms for agentic exploration.
Feature pgrp session cgroup PID ns branch()
Reliable termination
No escape possible
Sibling isolation
No setup required
No root privileges
No PID 1 complexity
Exclusive groups
Cgroup v2 delegation via systemd can avoid root but requires prior configuration.
3. Branch Contexts
We now introduce the branch context, a new OS abstraction that captures the requirements from Section 2. A branch context is an isolated execution environment, comprising a filesystem view and a process group, that provides copy-on-write state isolation (R1, R4), atomic commit with single-winner resolution (R2), nesting for hierarchical exploration (R3), lightweight unprivileged creation (R5), and coordinated process lifecycle management (R6). Figure 1 illustrates how the two components of our framework (BranchFS and branch()) realize this abstraction.
Parent Processbranch(N=3)Child 1Child 2Child 3Mount NSMount NSMount NSBranchFS (FUSE)Branch 1 ($\Delta$)Branch 2 ($\Delta$)Branch 3 ($\Delta$)Base Directory (Original Files)creates N childrenisolated viewscopy-on-write
Figure 1. Architecture overview. branch() coordinates process creation and mount namespace isolation; BranchFS provides filesystem branching with copy-on-write semantics. Dotted arrows indicate copy-on-write fallback to base.
3.1. Definition
A branch context $b_{i}$ encapsulates: (1) a filesystem view consisting of the origin’s files overlaid with a copy-on-write delta layer $\Delta_{i}$, and (2) a process group whose side effects are confined to $b_{i}$. Branch contexts are created in sets of $N$ siblings from a single frozen origin, execute independently, and resolve via atomic commit (promoting one context’s changes to the parent) or abort (discarding changes).
Prior work on speculative OS execution introduced related concepts: Tapus et al. (Tapus and Hickey, 2006) proposed “speculations” for distributed fault-tolerance, where a single speculative path runs sequentially from a checkpoint and an owner process decides whether to accept results. Branch contexts differ in three ways: they support parallel exploration of $N$ paths simultaneously, use competitive first-commit-wins resolution rather than owner-decided acceptance, and target local agent workloads rather than distributed system consistency.
3.2. Lifecycle: Fork, Explore, Commit
Every branch context follows a three-phase lifecycle:
Fork.
NN branch contexts are created atomically from a frozen origin. Each context $b_{i}$ receives its own copy-on-write filesystem layer and process group.
Explore.
Each branch context executes independently. Filesystem modifications are captured in the per-context delta layer $\Delta_{i}$ via copy-on-write; all process side effects are confined to the context.
Commit/Abort.
When a branch context succeeds (e.g., tests pass), it commits: its delta $\Delta_{i}$ is applied atomically to the parent. A context may also abort, discarding its delta with no effect on the parent or siblings.
3.3. Core Semantics
Branch contexts are defined by four core semantic properties:
  1. (1)
    Frozen origin. The parent state becomes read-only when branch contexts exist, eliminating merge conflicts by construction.
  2. (2)
    Parallel isolated execution. All $N$ contexts run simultaneously, not as sequential retries, and are fully isolated from one another: no context can observe or modify another’s state. This enables speedup proportional to available parallelism.
  3. (3)
    First-commit-wins. Any context may commit; the first to do so wins, and all siblings are invalidated. This avoids merge complexity and suits AI agent workloads where the orchestrator cannot predict which branch will succeed.
  4. (4)
    Nested branch contexts. A branch context may itself fork sub-contexts, forming a tree. Each level commits to its immediate parent, enabling hierarchical exploration (e.g., an agent tries three strategies, each of which tries two sub-variants).
The following sections describe how we realize branch contexts: BranchFS (Section 4) implements the filesystem dimension, and branch() (Section 5) implements the process coordination dimension.
4. BranchFS Design and Implementation
BranchFS is a FUSE-based filesystem that realizes the filesystem dimension of branch contexts.
4.1. File-Level Copy-on-Write
When a file is first modified on a branch, BranchFS copies the entire file from the base directory (or nearest ancestor branch that modified it) into the branch’s delta layer (implemented as a per-branch directory on disk). Subsequent reads and writes to that file are served entirely from the delta copy. Unmodified files are resolved by walking up the branch chain to the base directory.
This file-level granularity is a deliberate trade-off: it is coarser than block-level copy-on-write but simpler to implement correctly atop FUSE, and the overhead of copying whole files is acceptable for the small-to-medium files typical of agent workloads.
4.2. Branch Chain and Resolution
Branches form a chain from the current branch back through ancestor branches to the base directory. File lookups traverse this chain:
  1. (1)
    Check the current branch’s delta layer.
  2. (2)
    If not found, check each ancestor branch in order.
  3. (3)
    If not found in any branch, serve from the base directory.
  4. (4)
    If a tombstone marker is encountered at any level, the file is reported as deleted.
Tombstone markers are lightweight sentinel files that record deletions, ensuring deleted files do not “reappear” from the base directory.
4.3. Commit and Abort Semantics
Commit.
A commit applies a branch’s changes atomically to its parent: (1) Modified files and tombstones are collected from the committing branch’s delta layer. (2) Deletions (tombstones) are applied first. (3) Modified files are copied from the committing branch’s delta layer into the parent’s delta layer. (4) The parent’s epoch counter is incremented, invalidating all sibling branches. (5) Memory-mapped regions from invalidated branches trigger SIGBUS.
Abort.
An abort discards the branch’s delta layer without modifying the parent, incurring negligible cost. Sibling branches remain valid, and memory-mapped regions from the aborted branch trigger SIGBUS.
4.4. Mount-Based Isolation and Interface
BranchFS can be used standalone, without the branch() kernel extension. In standalone mode, branches are created, committed, and aborted via a CLI tool (branchfs create/commit/abort) or control files (.branchfs_ctl) within the filesystem, requiring no kernel modifications. All mounts of the same BranchFS daemon share a single branch namespace, with each branch accessible as a virtual directory via @branch paths (e.g., /mnt/workspace/@feature-a/). Per-branch isolation for multi-agent workflows is achieved by placing each agent in its own mount namespace with only its @branch path bind-mounted as the workspace root, preventing agents from accessing each other’s branches.
When used with branch() (Section 5), branch management is instead driven by the kernel: branch() issues FS_IOC_BRANCH_* ioctls to create, commit, and abort branches, and handles mount namespace setup atomically.
5. The branch() Syscall
While BranchFS provides filesystem branching, agents must still coordinate processes across branches. We propose branch(), a Linux syscall that provides this coordination with capabilities that are difficult or impossible to achieve purely in userspace: reliable process termination (no escape via setsid()), kernel-enforced sibling isolation, memory branching via page tables, and PID preservation on commit.
5.1. Interface
The branch() interface is designed around three principles: simplicity, exposing only three operations (create, commit, abort) that mirror the branching lifecycle; flexibility, allowing callers to select exactly which resources to branch via composable flags; and extensibility, using the bpf(2)-style multiplexed syscall pattern so new operations and flags can be added without changing the syscall signature.
The syscall takes an operation code and a union of per-operation argument structures (Listing 1).
Listing 1: The branch() syscall interface.
/* Operations */
#define BR_CREATE 1 /* Create branch(es) */
#define BR_COMMIT 2 /* Commit to parent */
#define BR_ABORT 3 /* Discard branch */
/* Flags for BR_CREATE */
#define BR_FS (1 << 0) /* Branch filesystem */
#define BR_MEMORY (1 << 1) /* Branch memory */
#define BR_ISOLATE (1 << 2) /* Sibling isolation */
#define BR_CLOSE_FDS (1 << 3) /* Close all fds */
/* Generic ioctls for branching filesystems */
#define FS_IOC_BRANCH_CREATE _IO(’b’, 0)
#define FS_IOC_BRANCH_COMMIT _IO(’b’, 1)
#define FS_IOC_BRANCH_ABORT _IO(’b’, 2)
union branch_attr {
struct { /* BR_CREATE */
__u32 flags;
__s32 mount_fd; /* workspace fd */
__u32 n_branches;
__u64 child_pids; /* pid_t __user *, output */
} create;
struct { /* BR_COMMIT */
__u32 flags; /* reserved */
} commit;
struct { /* BR_ABORT */
__u32 flags; /* reserved */
} abort;
long branch(int op, union branch_attr *attr,
size_t size);
Table 3 summarizes why kernel support is necessary for the core capabilities. Beyond individual capabilities, the fundamental argument for a syscall is atomic composition: setting up cgroups, PID namespaces, mount namespaces, filesystem branches, and signal barriers in userspace is a multi-step process with race windows between steps (e.g., a process can fork between cgroup creation and migration), fragile cleanup on partial failure, and error-prone ordering dependencies. branch() composes all of these atomically in a single call with kernel-side cleanup on failure, following the same rationale that motivated clone() over manual fork() + unshare() sequences.
Table 3. Justification for branch(). Strong justifications require kernel support; weak justifications are conveniences.
Capability Userspace? Justification
Atomic composition Impossible Strong
Memory branching Impossible Strong
Process termination Cgroups only Strong
Sibling isolation PID ns only Strong
Atomic mount setup User ns trick Weak
Parent read-only mprotect + FUSE Weak
First-commit-wins Atomics Weak
Multi-branch create Fork loop Weak
5.2. Operations and Flags
BR_CREATE.
Creates n_branches child processes, each in its own branch. The kernel issues FS_IOC_BRANCH_CREATE on mount_fd once per child to create filesystem branches, then atomically forks each child into a new mount namespace and bind-mounts the resulting branch fd over mount_fd using the new mount API (open_tree + move_mount). Returns 0 to the parent (with child_pids filled) and $1..N$ to each child (branch index). All branches created in a single call form an exclusive group: only one can commit successfully; others receive -ESTALE. Four flags control branch scope: BR_FS (required) gives each child a new mount namespace with its branch mounted as the workspace; BR_MEMORY additionally branches memory via page table copy-on-write, where writes fault into child-owned pages and the parent’s pages become read-only; BR_ISOLATE adds kernel-enforced signal/ptrace barriers between siblings via branch_id checks; BR_CLOSE_FDS closes inherited file descriptors so children re-open files in their branch context.
BR_COMMIT.
The kernel first wins the exclusive group race, then issues FS_IOC_BRANCH_COMMIT on the branch fd to apply filesystem changes, commits memory if BR_MEMORY was set, and terminates siblings (which receive -ESTALE on their next commit attempt). The committing child then optionally replaces the parent (analogous to execve()), taking over its PID so the caller sees a single continuous process.
BR_ABORT.
The kernel issues FS_IOC_BRANCH_ABORT on the branch fd to discard filesystem changes, then terminates the branch process. If all branches abort, the parent resumes.
Nested Branches.
Branches can create sub-branches by calling branch() again. A commit replaces only the immediate parent; further propagation requires separate commits.
5.3. Usage Example
Listing 2 shows the typical pattern for an AI agent exploring three approaches in parallel. The first branch to succeed commits its filesystem changes to the parent; siblings receive -ESTALE and exit.
Listing 2: Parallel exploration with first-commit-wins.
1int mnt = open("/mnt/workspace", O_PATH);
2pid_t pids[3];
3union branch_attr attr = {
4 .create = { .flags = BR_FS,
5 .mount_fd = mnt,
6 .n_branches = 3,
7 .child_pids = (uintptr_t)pids }
8};
9int idx = branch(BR_CREATE, &attr, sizeof(attr));
10if (idx == 0) {
11 /* Parent: wait for winner */
12 while (wait(NULL) > 0);
13 /* Winner’s changes now visible in base */
14} else {
15 /* Child: idx is 1, 2, or 3 */
16 bool ok = try_approach(idx);
17 if (ok) {
18 union branch_attr ca = {.commit = {0}};
19 int r = branch(BR_COMMIT, &ca, sizeof(ca));
20 if (r == -ESTALE) _exit(1); /* lost */
21 /* r == 0: FS + process committed atomically */
22 } else {
23 union branch_attr $a$ = {.abort = {0}};
24 branch(BR_ABORT, &$a$, sizeof($a$));
25 /* does not return */
26 }
27}
5.4. Parent and Child Semantics
Parent Read-Only.
While branches exist, the parent becomes read-only for both filesystem (BranchFS denies writes) and memory (pages marked read-only, writes return -EAGAIN). This eliminates conflicts: the parent spawns branches and waits; on commit, the winning branch replaces it.
Fork and Threads.
fork() within a branch is permitted: the child inherits the mount namespace (filesystem changes captured in delta) but has a separate address space (memory not part of branch state). Threads via clone(CLONE_VM) share both mount namespace and address space, so their changes are part of the branch. On commit or abort, all descendants are terminated; the kernel tracks them for reliable cleanup.
5.5. Filesystem-Agnostic Integration
branch() integrates with branching filesystems entirely through three generic ioctls (FS_IOC_BRANCH_CREATE, _COMMIT, _ABORT), following the pattern of existing generic ioctls such as FICLONE and FIEMAP. For FUSE-based filesystems like BranchFS, ioctls are forwarded to the daemon via FUSE_IOCTL; for kernel-native filesystems like DAXFS (Cong Wang, 2026), they are handled directly. Crucially, branch() itself contains no filesystem-specific code: adding a new branching filesystem requires only implementing the three ioctls, with no changes to branch() or the VFS layer.
6. Preliminary Evaluation
We evaluate BranchFS, our implemented FUSE filesystem. The branch() design is not yet implemented.
Experimental Setup.
All experiments run on a machine with an AMD Ryzen 5 5500U CPU (6 cores, 12 threads), 8 GB DDR4 RAM, and a 240 GB NVMe SSD (WDC WDS240G2G0C), running Pop!_OS 22.04 with Linux kernel 6.17. BranchFS is implemented in approximately 3,400 lines of Rust using the fuser library (FUSE 3 protocol). Each measurement is the median of 10 trials.
Branch Creation.
[…]
Table 4. Branch creation latency vs. base directory size.
[…]
100 292
1,000 317
10,000 310
Commit and Abort.
Commit overhead is proportional to modified data volume, not total filesystem size. For small modifications, commit completes in under 1 ms. Abort is similarly fast since it only removes the delta layer without copying files to the parent. Table 5 shows commit and abort latencies for various modification sizes.
Table 5. Commit and abort latency vs. modification size.
[…]
1 KB 317 315
100 KB 514 365
1 MB 2,100 890
I/O Throughput.
We measure sequential read and write throughput on a 50 MB file (64 KB blocks) in three configurations: native filesystem, BranchFS with regular FUSE, and BranchFS with FUSE passthrough. Table 6 summarizes the results.
Table 6. Sequential I/O throughput comparison (MB/s).
Mode Read Write
Native filesystem 8,800 576
BranchFS (regular FUSE) 1,655 631
BranchFS (passthrough) 7,236 719
With file descriptor caching, regular FUSE mode achieves 1,655 MB/s read throughput (19% of native); the gap reflects the FUSE kernel-to-userspace roundtrip. Write throughput slightly exceeds native because BranchFS treats fsync as a no-op for ephemeral branches (durability is enforced at commit time), avoiding the SSD sync cost. Passthrough mode (FOPEN_PASSTHROUGH) bypasses the daemon for unmodified files, reaching 7,236 MB/s read (82% of native). These numbers suffice for agent workloads dominated by LLM API latency (100 ms–10 s).
7. Related Work
Isolation Mechanisms.
Containers (Merkel, 2014) combine Linux namespaces (Kerrisk, 2013) with cgroups (Heo, 2015) for resource isolation, and gVisor (Google, 2024) adds kernel-level sandboxing. MVVM (Yang et al., 2024b) extends this model to cross-platform agent deployment via WebAssembly enclaves. At a finer granularity, lwCs (Litton et al., 2016) provide independent protection domains within a process for fast context switching, Dune (Belay et al., 2012) uses hardware virtualization for user-level sandboxing, and Capsicum (Watson et al., 2010) introduces capability-based sandboxing for Unix. These systems target general-purpose containment or intra-process isolation, whereas branch contexts focus on lightweight copy-on-write branching with coordinated commit/abort semantics across parallel exploration paths.
OS-Level Speculation and Transactions.
Speculator (Nightingale et al., 2005) pioneered speculative execution with output gating for distributed file systems, targeting network latency rather than agent exploration. TxOS (Porter et al., 2009) introduced OS-level ACID transactions, but transactions are flat, short-lived, and require deep kernel modifications. In contrast, branch contexts support nesting, persist for arbitrary durations, and operate entirely in userspace. Baumann et al. (Baumann et al., 2019) argue that fork() is no longer a good abstraction, and branch() shares this motivation, replacing implicit resource inheritance with explicit branching via composable flags.
8. Discussion and Future Work
Semantic Extensions.
Our current prototype isolates only filesystem state. External side effects (network, IPC, device I/O) are not rolled back on abort. A complete solution requires effect gating: buffering external actions until commit and discarding them on abort. Agent gateways such as Agentry (Wang, 2025), which mediate all agent-to-external communication through a central gateway, provide a natural interposition point, and Speculator (Nightingale et al., 2005) demonstrated output gating for distributed file systems. Extending these ideas to general-purpose I/O is future work. On the resolution side, BranchFS currently supports only single-winner semantics and cannot combine results from multiple branches. Future work will explore multi-branch merge, from file-level union of non-overlapping changes, through conflict detection for overlapping files, to semantic merge strategies.
Implementation Roadmap.
branch() is currently a design proposal. We plan to prototype it on Linux 6.19, initially supporting BR_FS and BR_ISOLATE, with memory branching (BR_MEMORY) as a follow-on effort due to the page table complexity involved. On the filesystem side, BranchFS’s file-level copy-on-write has known limitations: symlinks targeting absolute paths outside the branch resolve to unbranched files, hardlinks lose their link relationship on copy-on-write, special files (FIFOs, sockets, device nodes) are unsupported in delta layer, and disk space exhaustion causes -ENOSPC errors. These suffice for agent workloads but need addressing for general use.
Current agents (Claude Code, SWE-agent, OpenHands) primarily checkpoint via files rather than process memory, suggesting that BR_FS-only mode covers most current use cases. The BR_MEMORY flag provides future capability for agents with significant in-memory state (e.g., persistent Python interpreters or embedding caches). Integration requires minimal changes: the orchestrator creates branches and commits the winner, while agents run unmodified inside their branch’s mount namespace.
Although we focus on AI agents, the fork/explore/commit abstraction generalizes: with n_branches=1 it provides single-branch try-and-rollback useful for package management (try an upgrade, abort if it breaks) and system configuration tuning.
9. Conclusion
Agentic exploration requires isolating both filesystem and process state across parallel execution paths, with atomic commit and rollback semantics. We introduce the branch context abstraction to capture these requirements and present two complementary implementations in Linux. BranchFS is a FUSE-based filesystem that provides lightweight copy-on-write branching with O(1) creation and atomic commit-to-parent, all without root privileges. branch() is a proposed Linux syscall for kernel-enforced process coordination with reliable termination and first-commit-wins semantics, integrating with branching filesystems through generic FS_IOC_BRANCH_* ioctls that keep the syscall filesystem-agnostic. Preliminary evaluation shows O(1) branch creation and modification-proportional commit costs.

### v2

@@
 Device-mapper snapshots (Linux Kernel Documentation, 2023) require raw block devices with O(depth) read latency.
-DAXFS (Cong Wang, 2026) provides branching with commit and sibling invalidation but targets memory-backed storage and is not portable.
 At coarser granularity, VM snapshot systems (QEMU Project, 2024; VMware, 2021; Cully et al., 2008; Lagar-Cavilla et al., 2009) have whole-VM overhead; recent serverless snapshot work (Agache et al., 2020; Ustiugov et al., 2021; Du et al., 2020) reduces startup latency but does not address branching semantics; and CRIU (CRIU Project, 2024) requires full state serialization, all violating R5.
@@
 Table 1. Feature comparison of filesystem branching mechanisms.
-Feature OverlayFS/UnionFS Btrfs/ZFS DM-Snap DAXFS BranchFS
+Feature OverlayFS/UnionFS Btrfs/ZFS DM-Snap BranchFS
 Portable across FSes
@@
 When used with branch() (Section 5), branch management is instead driven by the kernel: branch() issues FS_IOC_BRANCH_* ioctls to create, commit, and abort branches, and handles mount namespace setup atomically.
+4.5. Agent Integration Library
+To bridge BranchFS with agent frameworks, we provide BranchContext (Wang and Zheng, 2026), a Python library that wraps BranchFS’s branching primitives into high-level exploration patterns.
+At its core, a Workspace context manager mounts BranchFS and exposes Branch objects with automatic commit-on-success and abort-on-failure semantics.
+On top of this, the library provides seven composable patterns that directly realize the agent strategies described in Section 2: Speculate races $N$ candidates and commits the first success; BestOfN runs $N$ candidates and commits the highest-scoring one; Reflexion retries sequentially, feeding failure feedback into the next attempt; TreeOfThoughts explores hierarchical strategy trees with nested branches; BeamSearch keeps the top-$K$ branches alive at each depth level; Tournament performs pairwise elimination via a judge function; and Cascaded starts with one attempt and adaptively fans out on failure.
+Each pattern manages branch lifecycle (creation, commit, abort) and process isolation (via fork(2) with optional timeout and resource limits) internally, so agent code only supplies the per-branch task logic.
 5. The branch() Syscall
@@
 branch() integrates with branching filesystems entirely through three generic ioctls (FS_IOC_BRANCH_CREATE, _COMMIT, _ABORT), following the pattern of existing generic ioctls such as FICLONE and FIEMAP.
-For FUSE-based filesystems like BranchFS, ioctls are forwarded to the daemon via FUSE_IOCTL; for kernel-native filesystems like DAXFS (Cong Wang, 2026), they are handled directly.
+For FUSE-based filesystems like BranchFS, ioctls are forwarded to the daemon via FUSE_IOCTL; for kernel-native filesystems, they are handled directly.
 Crucially, branch() itself contains no filesystem-specific code: adding a new branching filesystem requires only implementing the three ioctls, with no changes to branch() or the VFS layer.
@@
 The BR_MEMORY flag provides future capability for agents with significant in-memory state (e.g., persistent Python interpreters or embedding caches).
-Integration requires minimal changes: the orchestrator creates branches and commits the winner, while agents run unmodified inside their branch’s mount namespace.
+The BranchContext library (Section 4.5) already provides this integration for Python-based agents, wrapping BranchFS into ready-to-use exploration patterns so that agent code only supplies per-branch task logic.
+Agents run unmodified inside their branch’s mount namespace; the library manages branch lifecycle and process isolation.
 Although we focus on AI agents, the fork/explore/commit abstraction generalizes: with n_branches=1 it provides single-branch try-and-rollback useful for package management (try an upgrade, abort if it breaks) and system configuration tuning.
@@
 BranchFS is a FUSE-based filesystem that provides lightweight copy-on-write branching with O(1) creation and atomic commit-to-parent, all without root privileges.
+On top of BranchFS, the BranchContext Python library provides seven composable exploration patterns (parallel speculation, best-of-N, reflexion, tree-of-thoughts, and others) for direct integration with agent frameworks.
 branch() is a proposed Linux syscall for kernel-enforced process coordination with reliable termination and first-commit-wins semantics, integrating with branching filesystems through generic FS_IOC_BRANCH_* ioctls that keep the syscall filesystem-agnostic.